# Multilingual DistilWhisper: Efficient Distillation of Multi-Task Speech Models via Language-Specific Experts

### v3

@@ Abstract
Whisper is a multitask and multilingual speech model covering 99 languages. It yields commendable automatic speech recognition (ASR) results in a subset of its covered languages, but the model still underperforms on a non-negligible number of under-represented languages, a problem exacerbated in smaller model versions. In this work, we propose DistilWhisper, an approach able to bridge the performance gap in ASR for these languages while retaining the advantages of multitask and multilingual capabilities. Our approach involves two key strategies: lightweight modular ASR fine-tuning of whisper-small using language-specific experts, and knowledge distillation from whisper-large-v2. This dual approach allows us to effectively boost ASR performance while keeping the robustness inherited from the multitask and multilingual pre-training. Results demonstrate that our approach is more effective than standard fine-tuning or LoRA adapters, boosting performance in the targeted languages for both in- and out-of-domain test sets, while introducing only a negligible parameter overhead at inference.
Index Terms— knowledge distillation, multitask speech processing, automatic speech recognition, multilingual speech processing, language experts
@@ 1 Introduction
Whisper [1] is a popular multilingual and multitask speech model that is known for its robustness (i.e. invariant performance over different out-of-domain data) for automatic speech recognition (ASR) [2]. This model covers 99 languages, and jointly trains on ASR, speech translation (many-to-English), language identification, and voice activity detection tasks. The original paper points this multitask training as a reason for the observed robustness of the model to out-of-domain data: compared to the English wav2vec 2.0 model [3], Whisper performance seems to generalize better to unseen domains. Presented in many sizes (from tiny to large-v2), we note that there is an important gap in ASR performance between whisper-large-v2 (largest model) and whisper-small (second smallest model) on a large set of languages, including low-resource languages, but also many high- and mid-resource ones. Such phenomenon in NLP is often referred as curse of multilinguality [4, 5, 6], where the performance drop due to the growing amount of covered languages can only be recovered via extensive model scaling. Such scaling comes with an important inference cost increase: for instance, whisper-large-v2 is 2-3 times slower than whisper-small.
@@
A common approach to efficient inference is distilling knowledge from a large multilingual teacher model into a smaller model [7, 8]. However, to apply such knowledge distillation (KD) to whisper-large-v2, the best and largest Whisper model, we would need to access unavailable information such as the training data across all the tasks and languages, in order to preserve the robustness of the original model. Recent works [9, 10] have demonstrated that the curse of multilinguality can also be solved by equipping a moderately sized model with language-specific (LS) modules. Such architecture allows to extend model parameters via extra modules when more languages are added into the model, thus maintaining consistent performance across languages, with no (or very low) extra computations at inference.
@@
Inspired by those findings we propose DistilWhisper, which extends whisper-small with LS feed-forward layers, that are used in parallel with the original feed-forward layers of the model. In order to preserve the robustness of the original model, DistilWhisper introduces the following extensions of previous works: (1) Following [11] we extend conditional language-specific routing (CLSR) modules with the gating mechanism that can route input representation either through the original feed-forward layer or through newly learned LS feed-forward layer; (2) When learning LS layers, we use whisper-large-v2 as a teacher model with the hypothesis that the KD loss should help reproducing the robustness of the larger Whisper model.
@@
Through extensive experiments on a diverse set of languages we demonstrate the effectiveness of DistilWhisper compared to standard fine-tuning or LoRA [12] adapters. Our lightweight ASR fine-tuning approach based on CLSR modules generalizes better than LoRA, and the introduction of KD further boosts results in both in- and out-of-domain test sets. We perform additional ablation studies showing our approach can cope with different amounts of training data. Finally, we demonstrate that the flexibility introduced by the gating mechanism equips DistilWhisper with an efficient adaptation approach, leveraging the LS modules only when those are relevant. We make available the models’ weights111Weights available at: https://huggingface.co/collections/naver/multilingual-distilwhisper-6576ecae8d209fc6a767d9e7. and code222Code available at: https://github.com/naver/multilingual-distilwhisper developed in this work.
@@ 2 Background
State of the art for ASR: Current approaches for ASR mainly rely on the adaptation of pre-trained Transformer stacks learned through self-supervision (i.e. SSL models) on unlabeled audio data. Such pre-trained models vary on the usage of pretext tasks [3, 13, 14] and language coverage [15, 16, 10, 17]. In contrast to this branch of research, the Whisper model relies on weak supervision, which means that the architecture is trained on weakly labeled data only (no self-supervision). Nonetheless, they show that with sufficient amounts of data, the model reaches competitive results compared to mono [1, 2] and multilingual SSL models [10].
Knowledge distillation (KD) has been initially proposed by [18] to distill knowledge from ensemble of models into a single model for ASR. It has further been used to distill knowledge from a large teacher model into smaller student models [7, 8, 19]. While original KD methods relied on minimization of KL-divergence between a teacher model and a student model, recent research [20, 21] have shown that symmetric divergences, such as Jensen-Shannon (JS) divergence, suffer less from borderline behaviors and lead to better results on sequence level distillation.
Adapters are small lightweight modules which are commonly used to adapt pre-trained models to new tasks or domains. In speech-related tasks, adapter-based fine-tuning has been utilized for speech translation [22, 23, 24], and domain adaptation [25, 26], for which they exhibit a similar performance to standard fine-tuning, but with only a fraction of trainable parameters. We also find work on task-adaptation of Whisper [27, 28, 29] using LoRA adapters. In contrast to adapters, in this work we introduce gated LS layers into Whisper, and propose a parameter-efficient KD approach that allows us to increase robustness to out-of-domain data.
@@ 3 DistilWhisper
With the goal of increasing performance for different languages in models of limited capacity, we propose the DistilWhisper approach: we plug conditional language-specific routing (CLSR) modules [11] into a small Whisper (whisper-small), and optimize these modules jointly on ASR fine-tuning and KD from a larger Whisper (whisper-large-v2). Figure 1 presents our architecture, below we detail its key components.
@@
CLSR module. We extend CLSR modules for the first time to the speech domain. This module learns a hard binary gate g(·) for each input token by using its hidden embedding . These decisions enable a layer to selectively guide information through either a LS path denoted as or a shared path referred to as , as in Eq 1. In contrast to the original CLSR, in this work we use LS gates as shown in Figure 1, instead of sharing them across languages. This allows us to train LS components individually (i.e. in parallel), and then only load the relevant modules at inference. Moreover, our approach also differs from the original CLSR by the positioning: supported by previous work [11, 9], we limit CLSR to the feed-forward, which we also replace entirely by the CLSR module, reducing further the number of parameters. Gating follows [11]: each gate is made by a two-layer bottleneck network, which is summed to an increasing zero-mean Gaussian noise during training in order to discretize it. At inference time, we adopt hard gating.
@@
DistilWhisper approach is detailed at Figure 1. Our student is enriched with CLSR modules at each feed-forward for each language. These CLSR layers are initialized from the frozen weights of the corresponding feed-forward layer. At training time, for each language the model updates only the corresponding LS layers and gates. At inference time, the model loads the shared layers (multilingual) and the LS modules and gates for the languages of interest, resulting in a limited parameter overhead. We highlight that the use of CLSR modules brings more flexibility to our architecture when compared to adapters, as it allows for routing at the token-level. This makes this approach more capable of leveraging pre-existing knowledge (shared frozen module) via LS gating activation.
DistilWhisper optimization. Following [11], when learning CLSR module parameters, in addition to standard cross-entropy loss , we employ a gate budget loss (Eq 2) to balance models’ usage of LS and language-shared modules. It relies on the gate activation values for a pair (audio, text) in a batch , which is expressed by where and are respectively the encoders and decoders layers, and when LS layer is selected, or otherwise. The average of this gate usage is constrained to a budget (Eq 2). For KD, following recent research [20, 21], we use JS divergence, whose loss is detailed in Eq 3, where is the teacher distribution, is the student distribution, and are sampled from the teacher’s and student’s distributions and compared with their average . Thus, CLSR modules parameters are learned to minimize final loss expressed as .
@@ 4 Experimental Setup
Datasets: We downsample the train and validation sets of the CommonVoice 13.0 (CV-13) dataset [30], using equal amounts of training data for each selected language: 10k utterances for training (approx. 14 h), 1k for validation. Data selection depends on the amount of up-votes utterances received by annotators. We do not downsample the test set. The FLEURS [31] dataset is used for out-of-domain evaluation, as it provides both a good language overlap with CV-13, and an effective out-of-domain setting for ASR evaluation. For instance, average number of tokens per sample for CV-13 is 36, and 97 for FLEURS.
@@
Language Selection: We consider all Whisper languages with a WER gap of more than 11 between large and small models on CV-13. We then narrow this list considering: 1) minimum amount of utterances (10k); 2) overlap with the FLEURS dataset. The final list of languages is: Catalan (ca), Czech (cs), Galician (gl), Hungarian (hu), Polish (pl), Thai (th), Tamil (ta) and Ukranian (uk).333Although Arabic would also qualify considering our criteria, we find that the dialect from FLEURS differs from the ones present on CV-13. These languages encompass 5 language sub-families and vary widely in terms of coverage in the Whisper training set, spanning from 4,300 h (pl) to just 9 h (gl).
@@
Models: We compare our approach to both whisper-small (pre-trained student) and whisper-large-v2 (teacher) models, as well as two approaches of fine-tuning (FT) for the student: standard fine-tuning (all weights are updated), and LoRA adaptation on top of the feed-forward layer. Finally, we also investigate the impact of the CLSR layer without the use of KD (CLSR-FT), decoupling the effect of KD from the flexibility offered by the routing mechanism on the consequent robustness of the model.
@@
Implementation: We train all models using the Transformers library [32], and make use of whisper-small and whisper-large-v2 pre-trained weights from HuggingFace.444https://huggingface.co/openai/ All models are trained for epochs using learning rate with linear decay, one epoch warm-up, batch size , and label smoothing factor . For LoRA, we use the hyperparameters proposed by [28]. For CLSR training we set gate budget and skip-gate probability . For KD we employ JS divergence with temperature , weighted such as the learning objective is . We report normalized WER using the Whisper normalization with a slight modification to avoid splitting numbers and latin-scripted text into individual characters in languages that do not use space delimitation (th). In all cases, the best model is chosen based on WER on the down-sampled CV-13 validation set.
| FLEURS | CV-13 | FLEURS (out-of-domain) | CV-13 (in-domain for FT only) | ||||||||||||||||
| #params | avg | avg | ca | cs | gl | hu | pl | ta | th | uk | ca | cs | gl | hu | pl | ta | th | uk | |
| whisper-large-v2 | 1.5B | 12.5 | 14.9 | 5.6 | 14.3 | 16.6 | 17.9 | 5.9 | 19.3 | 12.2 | 8.1 | 16.9 | 14.4 | 18.9 | 18.7 | 8.0 | 17.3 | 9.2 | 15.5 |
| whisper-small | 244M | 28.3 | 31.4 | 14.6 | 40.4 | 32.7 | 43.0 | 16.7 | 36.0 | 22.8 | 20.5 | 30.1 | 38.4 | 35.5 | 45.6 | 18.6 | 30.0 | 20.3 | 32.3 |
| whisper-small+FT | 244M | 23.30.06 | 16.30.09 | 15.5 | 31.0 | 16.9 | 36.7 | 22.0 | 22.7 | 15.6 | 25.9 | 13.7 | 20.5 | 11.3 | 24.1 | 16.3 | 13.6 | 7.4 | 23.4 |
| whisper-small+LoRA-FT | 379M | 24.90.07 | 18.20.02 | 17.6 | 36.9 | 18.2 | 41.6 | 25.9 | 15.2 | 11.7 | 31.8 | 14.0 | 23.7 | 12.7 | 28.0 | 21.2 | 12.0 | 7.9 | 26.4 |
| whisper-small+CLSR-FT | 369M | 23.40.19 | 16.30.08 | 15.7 | 30.5 | 17.2 | 36.9 | 22.8 | 22.7 | 15.6 | 25.8 | 14.1 | 20.3 | 11.6 | 24.3 | 16.1 | 13.3 | 7.4 | 23.4 |
| DistilWhisper | 369M | 22.80.21 | 16.00.04 | 15.3 | 30.2 | 16.7 | 36.9 | 21.4 | 21.8 | 15.1 | 24.9 | 13.8 | 20.0 | 11.8 | 24.0 | 15.9 | 12.6 | 7.2 | 23.1 |

### v4

# ♣ Télécom Paris, Institut Polytechnique de Paris, France

@@ Abstract
Whisper is a multitask and multilingual speech model covering 99 languages.
It yields commendable automatic speech recognition (ASR) results in a subset of its covered languages, but the model still underperforms on a non-negligible number of under-represented languages, a problem exacerbated in smaller model versions.
In this work, we propose DistilWhisper, an approach able to bridge the performance gap in ASR for these languages while retaining the advantages of multitask and multilingual capabilities.
Our approach involves two key strategies: lightweight modular ASR fine-tuning of whisper-small using language-specific experts, and knowledge distillation from whisper-large-v2. This dual approach allows us to effectively boost ASR performance while keeping the robustness inherited from the multitask and multilingual pre-training.
Results demonstrate that our approach is more effective than standard fine-tuning or LoRA adapters, boosting performance in the targeted languages for both in- and out-of-domain test sets, while introducing only a negligible parameter overhead at inference.
Keywords: knowledge distillation, multitask speech processing, automatic speech recognition, multilingual speech processing, language experts
@@ 1 Introduction
Whisper (radford2023robust) is a popular multilingual and multitask speech model that is known for its robustness (i.e. invariant performance over different out-of-domain data) for automatic speech recognition (ASR) (gandhi2022esb). This model covers 99 languages, and jointly trains on ASR, speech translation (many-to-English), language identification, and voice activity detection tasks. The original paper attributes this multitask training as a reason for the observed robustness of the model to out-of-domain data: compared to the English wav2vec 2.0 model (baevski2020wav2vec), Whisper performance seems to generalize better to unseen domains. Available in many sizes (from tiny to large-v2), Whisper exhibits an important gap in ASR performance between whisper-large-v2 (largest model) and whisper-small (second smallest model) on a large set of languages, including low-resource languages, but also many high- and mid-resource ones. This phenomenon in NLP is often referred as curse of multilinguality, where the performance drop due to the growing amount of covered languages can only be recovered via extensive model scaling (mmmt_google; conneau2020unsupervised; goyal2021largerscale). Such scaling comes with an important inference cost increase: for instance, whisper-large-v2 is 2-3 times slower than whisper-small.
A common approach to efficient inference is distilling knowledge from a large multilingual teacher model into a smaller model (sanh2020distilbert; mohammadshahi-etal-2022-small). However, applying knowledge distillation (KD) to whisper-large-v2, the best and largest Whisper model, presents a challenge. We would ideally need access to training data across all tasks and languages to preserve robustness, but such data is unavailable, making it challenging to maintain the model’s out-of-domain generalization capabilities.
@@
In another direction, pfeiffer-etal-2022-lifting and pratap2023scaling have demonstrated that the curse of multilinguality can also be solved by equipping a moderately sized model with language-specific (LS) modules. Such architectures allow extending model parameters via extra modules when more languages are added into the model, thus maintaining consistent performance across languages, with no (or very low) extra computations at inference.
@@
Inspired by these findings, we propose DistilWhisper, which extends whisper-small with language-specific feed-forward modules, that are used in parallel with the original feed-forward layers of the model. In order to preserve the robustness of the original model, DistilWhisper introduces the following extensions of previous works:
1.
Following zhang2021share, we extend conditional language-specific routing (CLSR) modules with the gating mechanism that can route input representation either through the original feed-forward module or through newly learned LS feed-forward module;
2.
When learning language-specific modules, we use whisper-large-v2 as a teacher model with the hypothesis that the KD loss should help reproduce the robustness of the larger Whisper model.
@@
Through extensive experiments on a diverse set of languages we demonstrate the effectiveness of DistilWhisper compared to standard fine-tuning or LoRA adapters (hu2021lora). Our lightweight ASR fine-tuning approach based on CLSR modules generalizes better than LoRA, and the introduction of KD further boosts results in both in- and out-of-domain test sets. We perform additional ablation studies showing our approach can cope with different amounts of training data. Finally, we demonstrate that the flexibility introduced by the gating mechanism equips DistilWhisper with an efficient adaptation approach, leveraging the language-specific modules only when those are relevant. We make available the models’ weights111Weights available at: https://huggingface.co/collections/naver/multilingual-distilwhisper-6576ecae8d209fc6a767d9e7. and code222Code available at: https://github.com/naver/multilingual-distilwhisper. developed in this work.
@@ 2 Background
2.1 State of the art for ASR
Current approaches for ASR mainly rely on the adaptation of pre-trained Transformer stacks learned through self-supervision (i.e. SSL models) on unlabeled audio data. Such pre-trained models vary on the usage of pretext tasks (baevski2020wav2vec; hsu2021hubert; chen2022wavlm) and language coverage (conneau21_interspeech; babu22_interspeech; pratap2023scaling; zhang2023google). In contrast to this branch of research, the Whisper model relies on weak supervision, which means that the architecture is trained on weakly labeled data only (no self-supervision). Nonetheless, radford2023robust show that with sufficient amounts of data, the Weakly Supervised Whisper model reaches competitive results compared to monolingual and multilingual SSL models (gandhi2022esb; pratap2023scaling).
2.2 Knowledge distillation
Knowledge distillation (KD) has been initially proposed by kd_hinton to distill knowledge from ensemble of models into a single model for ASR. It has further been used to distill knowledge from a large teacher model into smaller student models (sanh2020distilbert; mohammadshahi-etal-2022-small; shen2023language). While original KD methods relied on minimization of KL-divergence between a teacher model and a student model, js_vs_kd_acl23 and js_vs_kd_bb2023 have recently shown that symmetric divergences, such as Jensen-Shannon (JS) divergence, suffer less from borderline behaviors and lead to better results on sequence level distillation.
2.3 Parameter-efficient Fine-tuning
Adapters are small lightweight modules which are commonly used in NLP to adapt pre-trained models to new tasks or domains. In speech-related tasks, adapter-based fine-tuning has been utilized for speech translation (le2021lightweight; gow2023naver; antonios2022findings), and domain adaptation (thomas2022efficient; tomanek2021residual), for which they exhibit a similar performance to standard fine-tuning, but with only a fraction of trainable parameters. We also find work on task-adaptation of Whisper (radhakrishnan2023parameter; wang23ga_interspeech; feng2023peft) using LoRA adapters. In contrast to adapters, in this work we introduce gated LS modules into Whisper, and propose a parameter-efficient KD approach that allows us to increase robustness to out-of-domain data.
@@ 3 DistilWhisper
With the goal of increasing performance for different languages in models of limited capacity, we propose the DistilWhisper approach: we plug conditional language-specific routing (CLSR) modules (zhang2021share) into a small Whisper (whisper-small), and optimize these modules jointly on ASR fine-tuning and KD from a larger Whisper (whisper-large-v2). Figure Multilingual DistilWhisper: Efficient Distillation of Multi-Task Speech Models via Language-Specific Experts presents our architecture, below we detail its key components.
3.1 Conditional Language-Specific Routing (CLSR) module
We extend CLSR modules for the first time to the speech domain. This module learns a hard binary gate g(·) for each input token by using its hidden embedding . These decisions enable a layer to selectively guide information through either a Language-Specific path denoted as or a shared path referred to as , as in Eq 1:
@@ 3.1 Conditional Language-Specific Routing (CLSR) module
In contrast to the original CLSR, in this work we use language-specific gates as shown in Figure Multilingual DistilWhisper: Efficient Distillation of Multi-Task Speech Models via Language-Specific Experts, instead of sharing them across languages. This allows us to train LS modules individually (i.e. in parallel), and then only load the relevant modules at inference. Moreover, our approach also differs from the original CLSR by the positioning: supported by zhang2021share and pfeiffer-etal-2022-lifting works, we limit CLSR to the feed-forward, which we entirely replace with the CLSR module, reducing further the number of parameters.
Following zhang2021share, the gating mechanism is implemented as follows: each gate consists of a two-layer bottleneck network, to which we add an increasing zero-mean Gaussian noise during training to facilitate discretization and enable gradient flow through the binary decisions. At inference time, we adopt hard gating, where the gate outputs are deterministically set to either 0 or 1 based on the learned routing decisions.
3.2 DistilWhisper approach
@@
DistilWhisper approach is detailed in Figure Multilingual DistilWhisper: Efficient Distillation of Multi-Task Speech Models via Language-Specific Experts. Our student is enriched with CLSR modules at each feed-forward for each language. These CLSR layers are initialized from the frozen weights of the corresponding feed-forward layer. At training time, for each language the model updates only the corresponding language-specific modules and gates. At inference time, the model loads the shared modules (multilingual) and the LS modules and gates for the languages of interest, resulting in a limited parameter overhead. We highlight that the use of CLSR modules brings more flexibility to our architecture when compared to adapters, as it allows for routing at the token-level. This makes this approach more capable of leveraging pre-existing knowledge (shared frozen module) via LS gating activation.
3.3 DistilWhisper optimization
Following zhang2021share, when learning CLSR module parameters, in addition to standard cross-entropy loss , we employ a gate budget loss (Eq 3) to balance models’ usage of LS and language-shared modules.
The gate budget loss relies on the gate activation values for a pair (audio, text) in a batch , which is expressed by:
where and are respectively the encoders and decoders layers, and when LS module is selected, or otherwise. The average of this gate usage is constrained to a budget (Eq 3):
For KD, following js_vs_kd_acl23 and js_vs_kd_bb2023, we use JS divergence, whose loss is detailed in Eq 4:
where is the teacher distribution, is the student distribution, and are sampled from the teacher’s and student’s distributions and compared with their average .
Thus, CLSR modules parameters are learned to minimize final loss expressed as:
@@ 4.1 Datasets
We downsample the train and validation sets of the CommonVoice 13.0 (CV-13) dataset (ardila-etal-2020-common), using equal amounts of training data for each selected language: 10k utterances for training (approx. 14 h), 1k for validation. Data selection depends on the amount of up-votes utterances received by annotators. We do not downsample the test set. The FLEURS (conneau2023fleurs) dataset is used for out-of-domain evaluation, as it provides both a good language overlap with CV-13, and an effective out-of-domain setting for ASR evaluation. For instance, average number of tokens per sample for CV-13 is 36, and 97 for FLEURS.
4.2 Language Selection
@@
We consider all Whisper languages with a WER gap of more than 11 between large and small models on CV-13. We then narrow this list considering: 1) minimum amount of utterances (10k); 2) overlap with the FLEURS dataset. The final list of languages is: Catalan (ca), Czech (cs), Galician (gl), Hungarian (hu), Polish (pl), Thai (th), Tamil (ta) and Ukrainian (uk).333Although Arabic would also qualify considering our criteria, we find that the dialect from FLEURS differs from the ones present on CV-13. These languages encompass 5 language sub-families and vary widely in terms of coverage in the available Whisper training data, spanning from 4,300 h (pl) to just 9 h (gl).
4.3 Models
@@
We compare our approach to both whisper-small (pre-trained student) and whisper-large-v2 (teacher) models, as well as two approaches of fine-tuning (FT) for the student: standard fine-tuning (all weights are updated), and LoRA adaptation on top of the feed-forward layer. Finally, we also investigate the impact of the CLSR layer without the use of KD (CLSR-FT), decoupling the effect of KD from the flexibility offered by the routing mechanism on the consequent robustness of the model.
4.4 Implementation details
@@
We train all models using the Transformers library (wolf2020transformers), using the pre-trained weights for whisper-small and whisper-large-v2 available on HuggingFace444Models weights are available at: https://huggingface.co/openai/whisper-small and https://huggingface.co/openai/whisper-large-v2. All models are trained for epochs using a learning rate of with linear decay, one epoch of warm-up, batch size of 16, and label smoothing factor of 0.1. For LoRA, we use the hyperparameters proposed by wang23ga_interspeech. For CLSR training, we set the gate budget and skip-gate probability . For knowledge distillation (KD), we employ the JS divergence with temperature . The full learning objective for our experiments is given by:
We report normalized WER using the Whisper normalization procedure, with a slight modification to avoid splitting numbers and latin-scripted text into individual characters for languages that do not use space delimitation (e.g., Thai). In all cases, the best model is chosen based on WER on the down-sampled CV-13 validation set.
| FLEURS (out-of-domain) | ||||||||||
| Model | # params | avg | ca | cs | gl | hu | pl | ta | th | uk |
| whisper large-v2 | 1.5 B | 12.5 | 5.6 | 14.3 | 16.6 | 17.9 | 5.9 | 19.3 | 12.2 | 8.1 |
| whisper-small | 244 M | 28.3 | 14.6 | 40.4 | 32.7 | 43.0 | 16.7 | 36.0 | 22.8 | 20.5 |
| whisper-small + FT | 244 M | 23.30.06 | 15.5 | 31.0 | 16.9 | 36.7 | 22.0 | 22.7 | 15.6 | 25.9 |
| whisper-small + LoRA-FT | 379 M | 24.90.07 | 17.6 | 36.9 | 18.2 | 41.6 | 25.9 | 15.2 | 11.7 | 31.8 |
| whisper-small + CLSR-FT | 369 M | 23.40.19 | 15.7 | 30.5 | 17.2 | 36.9 | 22.8 | 22.7 | 15.6 | 25.8 |
| DistilWhisper | 369 M | 22.80.21 | 15.3 | 30.2 | 16.7 | 36.9 | 21.4 | 21.8 | 15.1 | 24.9 |
| Common Voice 13.0 (in-domain for FT only) | ||||||||||
| Model | # params | avg | ca | cs | gl | hu | pl | ta | th | uk |
| whisper large-v2 | 1.5 B | 14.9 | 16.9 | 14.4 | 18.9 | 18.7 | 8.0 | 17.3 | 9.2 | 15.5 |
| whisper-small | 244 M | 31.4 | 30.1 | 38.4 | 35.5 | 45.6 | 18.6 | 30.0 | 20.3 | 32.3 |
| whisper-small + FT | 244 M | 16.30.09 | 13.7 | 20.5 | 11.3 | 24.1 | 16.3 | 13.6 | 7.4 | 23.4 |
| whisper-small + LoRA-FT | 379 M | 18.20.02 | 14.0 | 23.7 | 12.7 | 28.0 | 21.2 | 12.0 | 7.9 | 26.4 |
| whisper-small + CLSR-FT | 369 M | 16.30.08 | 14.1 | 20.3 | 11.6 | 24.3 | 16.1 | 13.3 | 7.4 | 23.4 |
| DistilWhisper | 369 M | 16.00.04 | 13.8 | 20.0 | 11.8 | 24.0 | 15.9 | 12.6 | 7.2 | 23.1 |